OpenStudy (anonymous):

One number is 7 times as large as another. The sum of their reciprocals is 72/7. Find the two numbers.

OpenStudy (fibonaccichick666):

Any ideas on how to translate this sentence into some eqs?

OpenStudy (anonymous):

nope. lol

OpenStudy (fibonaccichick666):

check out here http://www.austincc.edu/tutor/students/resources/Translating%20English%20Words.pdf

OpenStudy (anonymous):

I don't understand any of that lol sorry

OpenStudy (fibonaccichick666):

I don't see what is so funny, and where don't you understand?

OpenStudy (anonymous):

Turning words into numbers

OpenStudy (fibonaccichick666):

did you read the site I just gave you?

OpenStudy (anonymous):

Yes

OpenStudy (fibonaccichick666):

it's rather self explanatory

OpenStudy (anonymous):

Ugh, I hate these.

OpenStudy (fibonaccichick666):

thst isn't really helpful. It just holds you back to say that. Anyways find "One number is 7 times as large as another" or something similar on that page I linked to, then tell me what it is

OpenStudy (anonymous):

the first 1 i guess. x+7?

OpenStudy (fibonaccichick666):

"ten more than x" uses none of the same words as "7 times as large"

OpenStudy (fibonaccichick666):

look

OpenStudy (fibonaccichick666):

It doesn't sound like you actually even read the site at the current moment in time

OpenStudy (anonymous):

none of them say large though...that's throwing me off

OpenStudy (fibonaccichick666):

did you scroll down to where it tells you what words mean what functions?

OpenStudy (anonymous):

Yup

OpenStudy (fibonaccichick666):

what does times mean?

OpenStudy (anonymous):

multiply

OpenStudy (fibonaccichick666):

so, if something is 7 times larger than a number,

OpenStudy (fibonaccichick666):

let "a number"= x

OpenStudy (anonymous):

@Rohith15 , why u delete it?

OpenStudy (anonymous):

it was wrong i took 7 greater than the first number

OpenStudy (anonymous):

i dont think the sum is 72/7

OpenStudy (anonymous):

well, I just need a formula to work with. I can probably go off of that

OpenStudy (anonymous):

see you take the two numbers as x and 7x. Now take their individual reciprocals i.e, 1/x and 1/7x. Their sum is given. Equate both sides . Take LCM. You will get an equation. Solve it.

OpenStudy (anonymous):

1/x+1/7x=72/7?

OpenStudy (anonymous):

yep. you got it!!

OpenStudy (anonymous):

So, I just added the numerator & denominator...I just got lost again.

OpenStudy (anonymous):

See when i simplified i got x = 1/9 an the other number is 9. But they are not satisfying, So I thought the sum must be wrong

OpenStudy (anonymous):

The 1/9 s correct, but not the 9. I jut checked the answer

OpenStudy (anonymous):

@wolf1728 , can you help me, please?

OpenStudy (anonymous):

i got it!!! the answers are 1/9 and 7/9. Just take the reciprocals nd add nd u get 72/7

OpenStudy (anonymous):

Thank you !!!!